OpenStudy (anonymous):

Find the angle θ (if it exists) in the interval [0°, 90°) for which sin θ = cos θ.

zepdrix (zepdrix):

\[\large \sin \theta = \cos \theta\]Divide both sides by cosine,\[\large \frac{\sin \theta}{\cos \theta}=1\]Recall your identity for this one,\[\large \frac{\sin \theta}{\cos \theta}=\tan \theta \qquad \rightarrow \qquad \tan \theta = 1\]What special angle, when you take the tangent, gives you 1?

OpenStudy (anonymous):

45?

zepdrix (zepdrix):

Yesss very good \c:/

OpenStudy (anonymous):

thank you!!